OpenStudy (cutestgirlaround):

The sum of these fractions is between which two whole numbers? 4 over 5 + 6 2 over 5 A. 5 and 6 B. 6 and 7 C. 7 and 8 D. 8 and 9

OpenStudy (retireed):

4 over 5 + 6 2 over 5 This is unclear. Repost with parenthesis and maybe I can help you.

OpenStudy (cutestgirlaround):

please help[

OpenStudy (retireed):

Oh I see it now. 4/5 + 6 and 2/5 = ???

OpenStudy (cutestgirlaround):

ok \[\frac{ 4 }{ 5 }\] + \[6\frac{ 2 }{ 5}\]

OpenStudy (retireed):

Convert to decimals 4/5 = 0.8 2/5 = 0.4 0.8 + 6.4 = _______

OpenStudy (cutestgirlaround):

7.2

OpenStudy (retireed):

yes and it is between ??? A. 5 and 6 B. 6 and 7 C. 7 and 8 D. 8 and 9

OpenStudy (cutestgirlaround):

7 and 8 ???

OpenStudy (retireed):

Is 7.2 between 7 and 8?

OpenStudy (cutestgirlaround):

hello

OpenStudy (retireed):

Hey

OpenStudy (retireed):

Is 7.2 between 7 and 8?

OpenStudy (cutestgirlaround):

sorry my laptop was being slow yeah i think its between 7 and 8

OpenStudy (retireed):

I get that problem a lot. Very good, you know the answer Bye!